OpenStudy (anonymous):

Decompose (-2x-23)/(2x^2 - 9x -5) into partial fractions

OpenStudy (watchmath):

can you factor 2x^2-9x-5 ?

OpenStudy (anonymous):

I'm not good at factoring

OpenStudy (anonymous):

would it be (2x+1)(x-5)

OpenStudy (watchmath):

yes, good! So now you want to find A and B such that \[ \huge \frac{-2x-23}{(2x+1)(x-5)}=\frac{A}{2x+1}+\frac{B}{x-5} \]

OpenStudy (anonymous):

How do I do that?

OpenStudy (watchmath):

Multiply both sides by \(\big (2x+1)(x-5)\). What do we get?

OpenStudy (anonymous):

I'm sorry lol but I don't know, I'm confused. Multiply that by what?

OpenStudy (anonymous):

@watchmath

OpenStudy (watchmath):

multiply each side of the equation by (2x+1)(x-5)